OpenStudy (anonymous):

Decide a solution to a differential equation 3y' -15y=0 that satisfies y'(0)=2 Thank you.

OpenStudy (kc_kennylau):

auxiliary equation: 3r-15=0, r=5 general solution: y=a*e^(5x) y'(x)=5a*e^(5x) y'(0)=5a=2, a=0.4 therefore solution is y=0.4*e^(5x)

OpenStudy (fifciol):

or like this: \[3\frac{ dy }{ dx }=15y \Rightarrow \frac{dy }{ y }=5dx\] integrate both sides \[\int\limits \frac{ dy }{ y}=5\int\limits dx\] \[lny=5x+C\] exponentiate: \[y=Ce^{5x} \Rightarrow y'=5Ce^{5x}\] I think you can finish this out now

OpenStudy (anonymous):

you can solve like this also. 3y'-15y=0 ory'-5y=0 when x=0,y'=2 2-5y=0 y=2/5 again y' =5y or dy/dx=5y dy/y=5dx \[\int\limits \frac{ dy }{y }=5\int\limits dx\] \[\ln y =5x+c\] \[y= e ^{5x+c}=e ^{c}e ^{5x}=Ce ^{5x}\] when x=0 ,y=2/5 \[\frac{ 2 }{5 }=C e ^{0}=C\] \[y=\frac{ 2 }{5 }e ^{5x}=0.4 e ^{5x}\]

OpenStudy (anonymous):

thank you all for the reply